Given the cubic equation 2z^3-z^2+4z+p=0, peR( p is an element of the set of real numbers)
z=1+2i (i) Find the other solution (ii) Determine the value of P
1
Expert's answer
2020-11-13T14:03:03-0500

"\displaystyle

f(z) = 2z^3-z^2+4z+p = 0\\


\textsf{By the factor theorem,}\, f(1 + 2i) = 0\\

2(1 + 2i)³ - (1 + 2i)² + 4(1 + 2i) + p = 0\\


2(1 + 6i - 12 - 8i) - (1 + 4i - 4) + 4 + 8i + p = 0\\


2(-2i - 11) - (-3 + 4i) + 4 + 8i + p = 0\\

-4i - 22 + 3 - 4i + 4 + 8i + p = 0\\

 \therefore p = 15\\


\textsf{By the Fundamental theorem of Algebra}\\
\textsf{the conjugate of}\, 1 + 2i \, \textsf{is also a root of}\, f(z)\\

\therefore (z - (1 + 2i))(z - (1 - 2i)) = 0\\

z^2 - (1 + 2i + 1 - 2i)z + (1 + 4) = 0\\

z^2 - 2z + 5 = 0\\

\textsf{By Long division, we obtain the last factor}\\

 2z + 3\\
z^2 - 2z + 5|2z^3-z^2+4z+15\\

 -(2z^3 - 4z^2 + 10z)\\
 
 3z^2 - 6z + 15\\
 -(3z^2 - 6z + 15)
 = 0\\


\implies(z^2 - 2z + 5)(2z + 3) = 0\\

2z + 3 = 0 \implies z = -\frac{3}{2}\\


\therefore z = -\frac{3}{2}\, \textsf{is another solution.}\\

(i)\\
-\frac{3}{2}\\

(ii)\\
p = 15"